# TESSER: Transfer-Enhancing Adversarial Attacks from Vision Transformers via Spectral and Semantic Regularization

### v1

@@ 1 Introduction
Deep learning models, particularly Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have achieved state-of-the-art performance across a broad spectrum of computer vision tasks [2, 49, 22]. Despite this progress, these models remain highly vulnerable to adversarial examples – carefully crafted perturbations that are imperceptible to humans but cause misclassification [10, 12, 13, 14]. In safety-critical applications such as autonomous driving and medical imaging, this fragility raises significant security concerns.
@@
Several recent works, such as ATT [24] and TGR [48], have explored ViT-specific mechanisms for improving transferability by truncating or regularizing gradient flows. However, these approaches either use fixed gradient masks or overlook token-level semantics, leading to suboptimal alignment with transferable visual features. In this paper, we introduce TESSER (Transfer-Enhancing Semantic and Spectral Regularization) a novel adversarial attack framework specifically designed to improve black-box transferability from ViT-based models to a diverse set of architectures. TESSER integrates two complementary strategies:
Feature-Sensitive Gradient Scaling (FSGS): a token-level gradient modulation method that scales gradients based on token importance derived from intermediate embeddings. Inspired by recent findings correlating token activation magnitudes with semantic relevance [19, 41, 25], FSGS steers the attack toward semantically meaningful regions and away from background or non-informative tokens, enhancing cross-model generalization.
@@ 2 Related Work
Adversarial Attacks on CNNs and ViTs. Adversarial attacks are small, human-imperceptible perturbations intentionally added to input data to mislead deep learning models [10]. For Convolutional Neural Networks (CNNs), numerous gradient-based attacks have been proposed to improve transferability, including momentum-based methods [4], variance tuning [18], and gradient skipping techniques [40]. These methods aim to stabilize perturbation updates and avoid local optima in the input space. However, attack techniques designed for CNNs do not transfer well to Vision Transformers (ViTs), which have fundamentally different architectures and information flow patterns. Recent works have proposed ViT-specific attacks that exploit token structure and attention mechanisms [26, 38]. For example, Token Gradient Regularization (TGR) [48] modifies intermediate-layer gradients to reduce token-wise variance, improving transferability within ViT families.
@@
Regularizing gradients is an effective way to suppress model-specific patterns and improve cross-model generalization. In CNNs, methods like SGM [40] and BPA [42] aim to manipulate the gradient flow through skip connections or rectify distortions introduced by nonlinearities. Others have employed gradient variance reduction [18] and ensemble-based tuning [44]. Attacks based on feature information [37, 9] focus on disrupting salient internal representations. However, improperly guided feature-based attacks risk discarding useful information and reducing transferability. To mitigate this, neuron attribution methods [48] and attention map diversification [28] have been explored, particularly in ViTs. ATT [24] introduces hybrid token gradient truncation by weakening gradients in attention and QKV blocks across layers of a ViT model. It leverages empirical observations of gradient variance to suppress high-magnitude gradients associated with overfitting, thereby improving transferability. However, ATT applies static truncation and does not explicitly consider token-level semantic relevance, which may limit its effectiveness when generalizing across diverse architectures. In contrast, our method introduces Feature-Sensitive Gradient Scaling (FSGS), which adaptively reweights gradients at a token level based on feature norms. This allows us to preserve semantically important gradients while suppressing noisy or architecture-specific ones, achieving improved transferability across ViTs, hybrids, and CNNs.
@@
Input Diversity and Spectral Regularization. Input diversity has been widely adopted to improve adversarial transferability. DI-FGSM [43] applies random resizing and padding, while PatchOut [38] discards patch-wise perturbations to prevent overfitting. Recent self-paced extensions further refine patch discarding based on semantic guidance [24]. While these approaches diversify the spatial patterns of inputs, few works address the frequency structure of perturbations. Our method incorporates Spectral Smoothness Regularization (SSR) by applying differentiable Gaussian blur during optimization. SSR suppresses high-frequency noise and promotes smooth perturbation patterns that generalize better across model architectures, particularly important for CNNs and early ViT layers that rely on localized features. Importantly, input diversity is orthogonal to our method, and can potentially be combined with TESSER for further gains. We provide additional results and analysis combining input diversity with our framework in the Appendix C.
@@ 3.1 Preliminaries
Let $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ denote an input image with ground-truth label $y\in\{1,2,\dots,K\}$, and let $f(\cdot)$ be a deep neural network classifier. The goal of an untargeted adversarial attack is to generate a perturbation $\delta$ such that the perturbed input $\mathbf{x}^{\text{adv}}=\mathbf{x}+\delta$ is misclassified by the model, i.e., $f(\mathbf{x}^{\text{adv}})\neq y$, while ensuring that $\|\delta\|_{\infty}\leq\epsilon$. Unlike CNNs that process local image regions hierarchically, Vision Transformers [7] operate on a sequence of non-overlapping image patches. Given an input image $\mathbf{x}$, it is partitioned into $N=\frac{HW}{P^{2}}$ patches of size $P\times P$, each linearly projected to a $D$-dimensional embedding, resulting in tokens $\{\mathbf{z}_{1},\dots,\mathbf{z}_{N}\}\subset\mathbb{R}^{D}$. A learnable classification token $\mathbf{z}_{\text{cls}}$ is prepended, yielding a token sequence $\mathbf{Z}^{(0)}\in\mathbb{R}^{(N+1)\times D}$, which is enriched with positional encodings. ViTs consist of a stack of $L$ transformer blocks. Each block contains a Multi-Head Self-Attention (MHSA) module and a Multi-Layer Perceptron (MLP) module, connected via residual connections and layer normalization (LN).
@@ 3.2 Feature-Sensitive Gradient Scaling (FSGS)
To improve transferability, we propose Feature-Sensitive Gradient Scaling (FSGS), a fine-grained gradient modulation strategy that steers adversarial updates toward semantically relevant tokens while suppressing gradients associated with model-specific or noisy patterns. Unlike prior methods such as ATT [24] and TGR [48], which rely on fixed truncation or uniform regularization, FSGS leverages intermediate transformer features to dynamically adjust gradient flow on a per-token basis.
@@
Token-Level Importance Estimation. Given a token embedding matrix $\mathbf{Z}\in\mathbb{R}^{T\times D}$, we estimate the importance of token $i$ using the activation norm $\alpha_{i}=\|\mathbf{z}_{i}\|_{2}$, which serves as a proxy for semantic saliency. This assumption is supported by prior work in both NLP and vision [19, 41, 25], which shows that activation magnitudes often correlate with token informativeness or attention saliency. For instance, Kobayashi et al. [19] and Wang et al. [25] argue that vector norms contribute substantially to a token’s influence, while Wu et al. [41] highlight the role of transformed token magnitudes in ViT explanations. These scores are min-max normalized:
@@
$s_{i}=\gamma_{\text{base}}+\lambda(1-\hat{\alpha}_{i}),\quad\mathbf{g}_{i}^{\text{FSGS}}=s_{i}\cdot\mathbf{g}_{i}$
@@
Semantic Gradient Preservation. In Vision Transformers, each token contributes to the final prediction through self-attention, MLP transformations, and CLS-token pooling. The gradient of the loss with respect to a token’s embedding indicates how much that token influences the model’s output. FSGS leverages the activation norm of intermediate token embeddings as a proxy for semantic importance: Higher activation norms typically correspond to visually meaningful content such as object boundaries, textures, or salient regions [21, 41]. By preserving gradients associated with these high-importance tokens and suppressing those from less informative or background areas, FSGS steers adversarial perturbations toward features that are more likely to generalize across architectures. This selective modulation avoids wasting the attack budget on noisy or model-specific components, thereby improving transferability.
To validate this design intuition, we compute Grad-CAM [30] heatmaps using the incorrect (adversarial) predictions of the black-box model and analyze their spatial alignment with perturbations generated by FSGS. Interestingly, even though the model misclassifies the input, the resulting Grad-CAM activations still highlight semantically meaningful regions (e.g., object boundaries or foreground), suggesting that FSGS preserves gradients aligned with true semantic features. This supports our hypothesis that activation norm correlates with token importance and that FSGS steers perturbations toward generalizable, transferable regions. See Section 4.4 for qualitative visualizations and Appendix A for further analysis.
@@ 3.3 Spectral Smoothness Regularization (SSR)
In addition to feature-aware gradient modulation, we introduce Spectral Smoothness Regularization (SSR) to suppress high-frequency perturbation artifacts that often limit cross-architecture transferability. SSR operates by applying a differentiable Gaussian blur to the adversarial input at each iteration, effectively enforcing a low-pass filter on the evolving perturbation. Let $\delta$ denote the adversarial perturbation and $\mathcal{G}_{\sigma}(\cdot)$ the Gaussian blur operator with standard deviation $\sigma$. The input passed to the model becomes: $\mathbf{x}_{adv}^{\text{blur}}=\mathcal{G}_{\sigma}(\mathbf{x}+\delta)$. The intuition behind SSR is grounded in observations from both signal processing and adversarial transferability literature: perturbations dominated by high-frequency components tend to overfit to surrogate model-specific features and fail to generalize across architectures [35, 46]. By enforcing spectral smoothness, we bias the perturbation towards lower frequencies, which are more likely to align with perceptually salient and transferable features.
Unlike input diversity methods [43] that rely on stochastic resizing or padding to enhance transferability, SSR deterministically constrains the spectral characteristics of the perturbation itself. It is also fundamentally different from smoothing-based defenses, as the blur is applied during optimization rather than at test time, making it an attack-side regularizer. SSR is computationally efficient, introduces no additional parameters, and is compatible with any gradient-based attack pipeline. In our implementation, we integrate SSR into each PGD iteration following the perturbation update step. As shown in our experiments, SSR synergizes with FSGS to improve attack success rates, especially in black-box settings and when transferring across model families with divergent inductive biases.
@@ 3.4 Module-Wise Gradient Modulation
Vision Transformers differ from CNNs not only in architecture but also in how features and gradients evolve with depth. Prior studies [24, 47, 26] have shown that deeper transformer layers tend to encode more specialized, model-specific patterns (particularly in the attention maps) which can harm the transferability of adversarial perturbations. To address this, we introduce a Module-wise gradient modulation strategy that suppresses unstable gradients in deep attention layers and softly attenuates the gradient flow in all modules (Attention, QKV, MLP) based on their layer depth. Inspired by ATT [24], our approach consists of two key components:

### v2

@@ 1 Introduction
Deep learning models, particularly Convolutional Neural Networks (CNNs) and Vision Transformers (ViTs), have achieved state-of-the-art performance across a broad spectrum of computer vision tasks Carion et al. (2020); Zhu et al. (2021); Ma et al. (2022). Despite this progress, these models remain highly vulnerable to adversarial examples – carefully crafted perturbations that are imperceptible to humans but cause misclassification Goodfellow et al. (2014); Guesmi et al. (2023, 2024a, 2024b). In safety-critical applications such as autonomous driving and medical imaging, this fragility raises significant security concerns.
@@
Several recent works, such as ATT Ming et al. (2024) and TGR Zhang et al. (2023), have explored ViT-specific mechanisms for improving transferability by truncating or regularizing gradient flows. However, these approaches either use fixed gradient masks or overlook token-level semantics, leading to suboptimal alignment with transferable visual features. In this paper, we introduce TESSER (Transfer-Enhancing Semantic and Spectral Regularization) a novel adversarial attack framework specifically designed to improve black-box transferability from ViT-based models to a diverse set of architectures. TESSER integrates two complementary strategies:
Feature-Sensitive Gradient Scaling (FSGS): a token-level gradient modulation method that scales gradients based on token importance derived from intermediate embeddings. Inspired by recent findings correlating token activation magnitudes with semantic relevance Kobayashi et al. (2020); Wu et al. (2024); Modarressi et al. (2022a), FSGS steers the attack toward semantically meaningful regions and away from background or non-informative tokens, enhancing cross-model generalization.
@@ 2 Related Work
Adversarial Attacks on CNNs and ViTs. Adversarial attacks are small, human-imperceptible perturbations intentionally added to input data to mislead deep learning models Goodfellow et al. (2014). For Convolutional Neural Networks (CNNs), numerous gradient-based attacks have been proposed to improve transferability, including momentum-based methods Dong et al. (2018a), variance tuning Huang et al. (2019), and gradient skipping techniques Wu et al. (2020b). These methods aim to stabilize perturbation updates and avoid local optima in the input space. However, attack techniques designed for CNNs do not transfer well to Vision Transformers (ViTs), which have fundamentally different architectures and information flow patterns. Recent works have proposed ViT-specific attacks that exploit token structure and attention mechanisms Naseer et al. (2021); Wei et al. (2022). For example, Token Gradient Regularization (TGR) Zhang et al. (2023) modifies intermediate-layer gradients to reduce token-wise variance, improving transferability within ViT families.
@@
Regularizing gradients is an effective way to suppress model-specific patterns and improve cross-model generalization. In CNNs, methods like SGM Wu et al. (2020b) and BPA Xiaosen et al. (2023) aim to manipulate the gradient flow through skip connections or rectify distortions introduced by nonlinearities. Others have employed gradient variance reduction Huang et al. (2019) and ensemble-based tuning Xiong et al. (2022). Attacks based on feature information Wang et al. (2021); Ganeshan et al. (2019) focus on disrupting salient internal representations. However, improperly guided feature-based attacks risk discarding useful information and reducing transferability. To mitigate this, neuron attribution methods Zhang et al. (2023) and attention map diversification Ren et al. (2025) have been explored, particularly in ViTs. ATT Ming et al. (2024) introduces hybrid token gradient truncation by weakening gradients in attention and QKV blocks across layers of a ViT model. It leverages empirical observations of gradient variance to suppress high-magnitude gradients associated with overfitting, thereby improving transferability. However, ATT applies static truncation and does not explicitly consider token-level semantic relevance, which may limit its effectiveness when generalizing across diverse architectures. In contrast, our method introduces Feature-Sensitive Gradient Scaling (FSGS), which adaptively reweights gradients at a token level based on feature norms. This allows us to preserve semantically important gradients while suppressing noisy or architecture-specific ones, achieving improved transferability across ViTs, hybrids, and CNNs.
@@
Input Diversity and Spectral Regularization. Input diversity has been widely adopted to improve adversarial transferability. DI-FGSM Xie et al. (2019) applies random resizing and padding, while PatchOut Wei et al. (2022) discards patch-wise perturbations to prevent overfitting. Recent self-paced extensions further refine patch discarding based on semantic guidance Ming et al. (2024). While these approaches diversify the spatial patterns of inputs, few works address the frequency structure of perturbations. Our method incorporates Spectral Smoothness Regularization (SSR) by applying differentiable Gaussian blur during optimization. SSR suppresses high-frequency noise and promotes smooth perturbation patterns that generalize better across model architectures, particularly important for CNNs and early ViT layers that rely on localized features. Importantly, input diversity is orthogonal to our method, and can potentially be combined with TESSER for further gains. We provide additional results and analysis combining input diversity with our framework in the Appendix C.
@@ 3.1 Preliminaries
Let $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ denote an input image with ground-truth label $y\in\{1,2,\dots,K\}$, and let $f(\cdot)$ be a deep neural network classifier. The goal of an untargeted adversarial attack is to generate a perturbation $\delta$ such that the perturbed input $\mathbf{x}^{\text{adv}}=\mathbf{x}+\delta$ is misclassified by the model, i.e., $f(\mathbf{x}^{\text{adv}})\neq y$, while ensuring that $\|\delta\|_{\infty}\leq\epsilon$. Unlike CNNs that process local image regions hierarchically, Vision Transformers Dosovitskiy et al. (2021) operate on a sequence of non-overlapping image patches. Given an input image $\mathbf{x}$, it is partitioned into $N=\frac{HW}{P^{2}}$ patches of size $P\times P$, each linearly projected to a $D$-dimensional embedding, resulting in tokens $\{\mathbf{z}_{1},\dots,\mathbf{z}_{N}\}\subset\mathbb{R}^{D}$. A learnable classification token $\mathbf{z}_{\text{cls}}$ is prepended, yielding a token sequence $\mathbf{Z}^{(0)}\in\mathbb{R}^{(N+1)\times D}$, which is enriched with positional encodings. ViTs consist of a stack of $L$ transformer blocks. Each block contains a Multi-Head Self-Attention (MHSA) module and a Multi-Layer Perceptron (MLP) module, connected via residual connections and layer normalization (LN).
@@ 3.2 Feature-Sensitive Gradient Scaling (FSGS)
To improve transferability, we propose Feature-Sensitive Gradient Scaling (FSGS), a fine-grained gradient modulation strategy that steers adversarial updates toward semantically relevant tokens while suppressing gradients associated with model-specific or noisy patterns. Unlike prior methods such as ATT Ming et al. (2024) and TGR Zhang et al. (2023), which rely on fixed truncation or uniform regularization, FSGS leverages intermediate transformer features to dynamically adjust gradient flow on a per-token basis.
@@
Why Does Token Activation Norm Reflect Saliency? In Vision Transformers, token activation norm has been empirically associated with semantic saliency. Prior works—including Kobayashi et al. Kobayashi et al. (2020), Modarressi et al. Modarressi et al. (2022b), and Wu et al. Wu et al. (2024)—demonstrate that tokens with higher embedding norms tend to correspond to class-relevant features and foreground objects. These studies show that token norms correlate with saliency and attention relevance, and can serve as a reliable proxy for identifying object-centric or discriminative regions. Our own Grad-CAM visualizations (Section 4.4) reinforce this observation: tokens with high activation norm consistently align with semantically meaningful image regions, such as objects or salient parts. This empirical alignment justifies using token norm as a saliency prior for guiding gradient modulation in our FSGS. From a theoretical perspective (see Appendix A), we argue that if high-norm tokens represent salient features that are reused across models, then modulating gradients based on token norm enhances the alignment between surrogate and target model gradients. This, in turn, increases the likelihood that the generated perturbations will transfer effectively across architectures.
Does alignment correlate with better transfer? Our assumption is that alignment between gradients and semantically meaningful features is critical for enhancing attack effectiveness and transferability. This is based on both theoretical rationale and empirical observations. Specifically, in our FSGS design, token importance is treated as a proxy for semantic relevance. Tokens with high importance often correspond to salient regions such as foreground objects or class-discriminative parts. Amplifying gradients from these tokens strengthens perturbations that disrupt the model’s core semantic understanding, leading to higher attack success and better cross-model transferability. This idea is rooted in the observation that semantically aligned perturbations are more likely to generalize across architectures, which often differ in their treatment of early, low-level features but converge on similar high-level semantics. As noted by Goodfellow et al. Goodfellow et al. (2014), adversarial vulnerability is closely tied to the alignment of gradients with semantically relevant directions. While originally discussed in the context of CNNs, this principle extends to Vision Transformers, where deeper layers increasingly represent abstract, transferable concepts Raghu et al. (2021); Bhojanapalli et al. (2021).
FSGS operationalizes this principle by scaling gradients according to token importance scores, thus enforcing alignment between the perturbation direction and the image’s semantic structure.
Why Feature-Sensitive Gradient Scaling (FSGS)? The goal of FSGS is to amplify gradients associated with semantically important tokens (typically those with high activations norms) and suppress less informative gradients that contribute less to the model’s prediction. However, in ViTs, early layers tend to capture low-level, architecture-dependent features (e.g., local textures, positional cues), which are not semantically aligned or transferable across models. Hence, using the raw importance score $(\alpha)$ to scale gradients in these layers would unintentionally amplify gradients from less transferable features. To address this, we adopt a dual-stage strategy: in early layers, we scale gradients using $(1-\alpha)$ to suppress noisy or low-level gradients, while in deeper, semantically richer layers, we use $\alpha$ to enhance gradients aligned with class-discriminative feature. This approach is motivated by established findings in the literature. Raghu et al. Raghu et al. (2021) note that early ViT layers extract general representations that diverge from CNNs, reducing transferability. Bhojanapalli et al. Bhojanapalli et al. (2021) highlight that semantic robustness correlates with deeper layers, whereas early representations are more noisy and less class-relevant. Kim et al. Kim et al. (2024) further observe that early-layer attention maps are more architecture sensitive and less effective for cross-model transfer.
Token-Level Importance Estimation. Given a token embedding matrix $\mathbf{Z}\in\mathbb{R}^{T\times D}$, we estimate the importance of token $i$ using the activation norm $\alpha_{i}=\|\mathbf{z}_{i}\|_{2}$, which serves as a proxy for semantic saliency. This assumption is supported by prior work in both NLP and vision Kobayashi et al. (2020); Wu et al. (2024); Modarressi et al. (2022a), which shows that activation magnitudes often correlate with token informativeness or attention saliency. For instance, Kobayashi et al. Kobayashi et al. (2020) and Wang et al. Modarressi et al. (2022a) argue that vector norms contribute substantially to a token’s influence, while Wu et al. Wu et al. (2024) highlight the role of transformed token magnitudes in ViT explanations. These scores are min-max normalized:
@@
Let $l\in\{1,\dots,L\}$ denote the index of the current transformer block, and let $\mathcal{E}\subset\{1,\dots,L\}$ be the set of early layers (e.g., $\mathcal{E}=\{1,\dots,k\}$). Define an indicator function:
$\beta^{(l)}=\begin{cases}1&\text{if }l\in\mathcal{E}\quad\text{(early layer)}\\ 0&\text{otherwise}\end{cases}$ (1)
The final scaling factor for token $i$ at layer $l$ is then computed as: $s_{i}^{(l)}=\gamma_{\text{base}}+\lambda\cdot\left[(1-\beta^{(l)})\cdot\hat{\alpha}_{i}+\beta^{(l)}\cdot(1-\hat{\alpha}_{i})\right]$. And the FSGS-modulated gradient is: $\mathbf{g}_{i}^{(l),\text{FSGS}}=s_{i}^{(l)}\cdot\mathbf{g}_{i}^{(l)}$
@@
Semantic Gradient Preservation. In Vision Transformers, each token contributes to the final prediction through self-attention, MLP transformations, and CLS-token pooling. The gradient of the loss with respect to a token’s embedding indicates how much that token influences the model’s output. FSGS leverages the activation norm of intermediate token embeddings as a proxy for semantic importance: Higher activation norms typically correspond to visually meaningful content such as object boundaries, textures, or salient regions Long et al. (2023); Wu et al. (2024). By preserving gradients associated with these high-importance tokens and suppressing those from less informative or background areas, FSGS steers adversarial perturbations toward features that are more likely to generalize across architectures. This selective modulation avoids wasting the attack budget on noisy or model-specific components, thereby improving transferability.
To validate this design intuition, we compute Grad-CAM Selvaraju et al. (2017) heatmaps using the incorrect (adversarial) predictions of the black-box model and analyze their spatial alignment with perturbations generated by FSGS. Interestingly, even though the model misclassifies the input, the resulting Grad-CAM activations still highlight semantically meaningful regions (e.g., object boundaries or foreground), suggesting that FSGS preserves gradients aligned with true semantic features. This supports our hypothesis that activation norm correlates with token importance and that FSGS steers perturbations toward generalizable, transferable regions. See Section 4.4 for qualitative visualizations and Appendix A for further analysis.
@@ 3.3 Spectral Smoothness Regularization (SSR)
In addition to feature-aware gradient modulation, we introduce Spectral Smoothness Regularization (SSR) to suppress high-frequency perturbation artifacts that often limit cross-architecture transferability. SSR operates by applying a differentiable Gaussian blur to the adversarial input at each iteration, effectively enforcing a low-pass filter on the evolving perturbation. Let $\delta$ denote the adversarial perturbation and $\mathcal{G}_{\sigma}(\cdot)$ the Gaussian blur operator with standard deviation $\sigma$. The input passed to the model becomes: $\mathbf{x}_{adv}^{\text{blur}}=\mathcal{G}_{\sigma}(\mathbf{x}+\delta)$. The intuition behind SSR is grounded in observations from both signal processing and adversarial transferability literature: perturbations dominated by high-frequency components tend to overfit to surrogate model-specific features and fail to generalize across architectures Tsipras et al. (2018); Yin et al. (2019). By enforcing spectral smoothness, we bias the perturbation towards lower frequencies, which are more likely to align with perceptually salient and transferable features.
Unlike input diversity methods Xie et al. (2019) that rely on stochastic resizing or padding to enhance transferability, SSR deterministically constrains the spectral characteristics of the perturbation itself. It is also fundamentally different from smoothing-based defenses, as the blur is applied during optimization rather than at test time, making it an attack-side regularizer. SSR is computationally efficient, introduces no additional parameters, and is compatible with any gradient-based attack pipeline. In our implementation, we integrate SSR into each PGD iteration following the perturbation update step. As shown in our experiments, SSR synergizes with FSGS to improve attack success rates, especially in black-box settings and when transferring across model families with divergent inductive biases.
@@ 3.4 Module-Wise Gradient Modulation
Vision Transformers differ from CNNs not only in architecture but also in how features and gradients evolve with depth. Prior studies Ming et al. (2024); Yosinski et al. (2014); Naseer et al. (2021) have shown that deeper transformer layers tend to encode more specialized, model-specific patterns (particularly in the attention maps) which can harm the transferability of adversarial perturbations. To address this, we introduce a Module-wise gradient modulation strategy that suppresses unstable gradients in deep attention layers and softly attenuates the gradient flow in all modules (Attention, QKV, MLP) based on their layer depth. Inspired by ATT Ming et al. (2024), our approach consists of two key components: